# Kullback–Leibler Potential for Non-Ergodic Replication Dynamics: An Information-Theoretic Second Law

### v1

@@ III.1 Image replication via Gaussian convolution
We model a single replication step as a smoothing (information loss) Markov map $T_{\sigma}$ induced by discrete Gaussian convolution in an image $I:\mathbb{Z}^{2}\to[0,1]$. Let the normalized Gaussian kernel be
$G_{\sigma}(u,v)\ :=\ \frac{1}{2\pi\sigma^{2}}\,\exp\!\Big{(}-\frac{u^{2}+v^{2}}{2\sigma^{2}}\Big{)},\qquad(u,v)\in\mathbb{Z}^{2},\ \sigma>0,$ (22)
and define the replication map by discrete convolution;
$(I\ast G_{\sigma})(x,y)\ :=\ \sum_{u,v\in\mathbb{Z}}I(x-u,y-v)\,G_{\sigma}(u,v).$ (23)
Let $p_{n}$ denote the (normalized) histogram of pixel intensities before step $n$, viewed as a probability distribution on a finite intensity alphabet $\mathcal{A}$; the one-step channel $T_{\sigma}$ induced by (23) yields
$q_{n}\ =\ p_{n}\,T_{\sigma},\qquad p_{n+1}\ =\ q_{n}.$ (24)
Ergodic vs. non-ergodic (blockwise) blur.
To model non-ergodicity, partition the image grid into disjoint blocks $\{\mathcal{B}_{j}\}_{j=1}^{m}$ and restrict the kernel support so that no mass crosses block boundaries. For $(x,y)\in\mathcal{B}_{j}$,
$(I\ast_{j}G_{\sigma})(x,y)\ :=\ \sum_{\begin{subarray}{c}u,v\in\mathbb{Z}\\ (x-u,y-v)\in\mathcal{B}_{j}\end{subarray}}I(x-u,y-v)\,G_{\sigma}(u,v),$ (25)
which induces a block diagonal channel $T_{\sigma}=\mathrm{diag}\bigl{(}(T_{\sigma})_{1},\dots,(T_{\sigma})_{m}\bigr{)}$ and thus satisfies the block-invariance condition (5). The “ergodic” case corresponds to using (23) globally (no blocking).
@@
At each step we compute (nats)
$H(q_{n})\$$:=\ -\sum_{a\in\mathcal{A}}q_{n}(a)\,\log q_{n}(a),$ (26)
$H_{\times}(n)\$$:=\ H(p_{n},q_{n})\ =\ -\sum_{a\in\mathcal{A}}p_{n}(a)\,\log q_{n}(a),$ (27)
$D_{\mathrm{KL}}(p_{n}\|q_{n})\$$:=\ \sum_{a\in\mathcal{A}}p_{n}(a)\,\log\frac{p_{n}(a)}{q_{n}(a)},$ (28)
$V(p_{n})\$$:=\ \inf_{\pi\in\Pi(p_{0})}D_{\mathrm{KL}}(p_{n}\|\pi),$ (29)
where $\Pi(p_{0})$ is the reachable steady set defined in (9). By Theorem 1, $V(p_{n})$ is nonincreasing along (24) in the non-ergodic (blockwise) case; empirically we also observe monotone $V$ in the ergodic case where the steady set is a singleton.
Figure 1: Ergodic (global) vs. non-ergodic (blockwise) Gaussian blurring on a $256\times 256$ image with $\sigma=1.5$ over $n_{\mathrm{steps}}=50$. Metrics $H$ (26), $H_{\times}$ (27), $D_{\mathrm{KL}}$ (28), and $V$ (29) are reported at each step. In the blockwise case, monotonicity of $V$ follows from Theorem 1.
III.2 DNA replication as a block-diagonal substitution process
@@
Single-molecule and theory studies show that DNA polymerases advance stochastically with forward/backward steps, proofreading, and context-dependent kinetics [10, 12, 11, 8, 9]. To capture non-ergodicity induced by compositional structure (e.g. AT- vs. GC-rich segments), we model replication as a block-diagonal Markov process on nucleotides.
@@
Let $\mathcal{X}=\{\mathrm{A},\mathrm{T},\mathrm{C},\mathrm{G}\}$ with blocks $\mathcal{X}_{1}=\{\mathrm{A},\mathrm{T}\}$ and $\mathcal{X}_{2}=\{\mathrm{C},\mathrm{G}\}$. A one-step substitution kernel (copying) with block invariance (5) is
@@
KLD potential under unique block invariants.
If each block chain has a unique invariant $\pi_{j}^{\ast}\in\mathcal{I}_{j}$, then by (11) and (15)
$V(p_{n})\ =\ \sum_{j=1}^{2}w_{j}(p_{0})\,D_{\mathrm{KL}}\!\bigl{(}p_{n}^{(j)}\ \|\ \pi_{j}^{\ast}\bigr{)},$ (34)
and Theorem 1 ensures $V(p_{n+1})\leq V(p_{n})$ and $V(p_{n})\to 0$ as $n\to\infty$.
@@
A simple coarse-grained link between proofreading bias and a chemical affinity $\Delta\mu$ reads
$\frac{\alpha^{\prime}}{\beta^{\prime}}\ \approx\ \frac{\alpha}{\beta}\,e^{\Delta\mu/k_{\mathrm{B}}T},\qquad\frac{\gamma^{\prime}}{\delta^{\prime}}\ \approx\ \frac{\gamma}{\delta}\,e^{\Delta\mu/k_{\mathrm{B}}T},$ (35)
suggesting that a stronger driving sharpens $\pi_{j}^{\ast}$ and accelerates the decay of $V$ (cf. kinetic proofreading intuition [5]).
@@
At each step, we record (in nats): $H(q_{n})$, $H_{\times}(n)=H(p_{n},q_{n})$, $D_{\mathrm{KL}}(p_{n}\|q_{n})$, and $V(p_{n})$.
with monotone $V$ ensured by Theorem 1.
Block-level coarse graining (optional).
Define AT/GC block error rates at step $n$ by
$e_{\mathrm{AT}}(n)\$$:=\ \sum_{x\in\{\mathrm{A},\mathrm{T}\}}\frac{p_{n}(x)}{w_{1}(p_{n})}\sum_{y\in\{\mathrm{C},\mathrm{G}\}}\widetilde{T}_{y|x},$ (36)
$e_{\mathrm{GC}}(n)\$$:=\ \sum_{x\in\{\mathrm{C},\mathrm{G}\}}\frac{p_{n}(x)}{w_{2}(p_{n})}\sum_{y\in\{\mathrm{A},\mathrm{T}\}}\widetilde{T}_{y|x}.$ (37)
The channel induced $2\times 2$ on the coarse alphabet $\{\mathrm{AT},\mathrm{GC}\}$ is
$P(\tilde{y}\mid\tilde{x})\ =\begin{pmatrix}1-e_{\mathrm{AT}}&e_{\mathrm{AT}}\\[2.0pt] e_{\mathrm{GC}}&1-e_{\mathrm{GC}}\end{pmatrix},\qquad\tilde{x},\tilde{y}\in\{\mathrm{AT},\mathrm{GC}\},$ (38)
with input weights $\mathbf{w}(n)=(w_{1}(p_{n}),w_{2}(p_{n}))$. The block-level mutual information (bits) is
$I(\tilde{X};\tilde{Y})\ =\ \sum_{\tilde{x},\tilde{y}}w_{\tilde{x}}(n)\,P(\tilde{y}\mid\tilde{x})\,\log_{2}\!\frac{P(\tilde{y}\mid\tilde{x})}{\sum_{\tilde{x}^{\prime}}w_{\tilde{x}^{\prime}}(n)\,P(\tilde{y}\mid\tilde{x}^{\prime})}$. (39)
Under the symmetric approximation $w_{1}=w_{2}=\tfrac{1}{2}$ and $e_{\mathrm{AT}}=e_{\mathrm{GC}}=\varepsilon$,

### v2

@@ 2.1 Image replication via Gaussian convolution
We model a single replication step as a Markov smoothing map $T_{\sigma}$ implemented by a discrete Gaussian convolution on $I:\mathbb{Z}^{2}\to[0,1]$ followed by re-quantization. The kernel coincides with $T_{\sigma}$ here. The kernel (discrete, renormalized) is given by:
$T_{\sigma}(u,v)\;:=\;\frac{\exp\!\bigl(-(u^{2}+v^{2})/(2\sigma^{2})\bigr)}{\displaystyle\sum_{a,b\in\mathbb{Z}}\exp\!\bigl(-(a^{2}+b^{2})/(2\sigma^{2})\bigr)},\qquad(u,v)\in\mathbb{Z}^{2},\ \sigma>0,$ (20)
so that $\sum_{u,v}T_{\sigma}(u,v)=1$. The (global) discrete convolution is
$(I\ast T_{\sigma})(x,y)\ :=\ \sum_{u,v\in\mathbb{Z}}I(x-u,y-v)\,T_{\sigma}(u,v),$ (21)
and in the non-ergodic (blockwise) variant the sum is restricted to pixels belonging to the same predefined region, which simply makes the $T_{\sigma}$ block diagonal.
Let $p_{n}$ denote the (normalized) mosaic block distribution, histogram, before step $n$. The convolution (21) followed by the same binning induces the one-step Markov channel in the mosaic block distribution, denoted $T_{\sigma}$ so that
$q_{n}\ =\ p_{n}\,T_{\sigma},\qquad p_{n+1}\ =\ q_{n},$ (22)
with row vectors acting on the right.
@@
We view the histogram on a finite state space $\mathcal{X}$ (intensity bins).
$H(q_{n})$ $=-\sum_{x\in\mathcal{X}}q_{n}(x)\,\log q_{n}(x),$ (23)
$H_{\times}(n)$ $=-\sum_{x\in\mathcal{X}}p_{n}(x)\,\log q_{n}(x),$ (24)
$D_{\mathrm{KL}}(p_{n}\|q_{n})$ $=\sum_{x\in\mathcal{X}}p_{n}(x)\,\log\frac{p_{n}(x)}{q_{n}(x)},$ (25)
$V(p_{n})$ $=\inf_{\pi\in\Pi(p_{0})}D_{\mathrm{KL}}(p_{n}\|\pi).$ (26)
For Fig. 2 (Appendix B), values are shown in bits (base 2). Each panel annotates a certified $\varepsilon$ and the corresponding $\delta(T_{\sigma})$ for its (blockwise) kernel.
Changing minimizers.
When a block admits multiple invariant measures ($\mathcal{I}_{j}$ not a singleton, e.g. reducible blocks), the minimizer $\pi^{\star}\!\in\!\Pi(p_{0})$ of $V(p_{n})$ can change along the trajectory even though $V$ remains nonincreasing. A concrete three-state reducible example with an explicit switching of the KLD minimizer is given in Appendix G.
[Uncaptioned image][Uncaptioned image]
Doeblin $\varepsilon=0.073$, Dobrushin $\delta=0.83$.
Figure 1:
Refer to captionRefer to caption
Doeblin $\varepsilon=0.051$, Dobrushin $\delta=0.88$.
Refer to captionRefer to caption
Doeblin $\varepsilon=0.019$, Dobrushin $\delta=0.93$.
Figure 2: Ergodic (global) vs. non-ergodic (blockwise) Gaussian blurring of a $256\times 256$ image ($\sigma=1.5$, $n_{\mathrm{steps}}=50$). At each step we report $H$ (23), $H_{\times}$ (24), $D_{\mathrm{KL}}$ (25), and $V$ (26). In the blockwise case, monotonicity of $V$ follows from Theorem 1. (Upper) $2\times 2$ blocks, (Middle) $4\times 4$ blocks, (Lower) $128\times 128$ blocks. All values in this figure are in bits (base 2; see Appendix B). Per-panel mixing strength (see Defs. (7)–(8)). Upper: $\varepsilon=0.073$, $\delta=0.83$;   Middle: $\varepsilon=0.051$, $\delta=0.88$;   Lower: $\varepsilon=0.019$, $\delta=0.93$.
2.2 DNA replication as a block-diagonal substitution process
@@
Second, we model DNA replication with the proofreading/repairing mechanism after copying. DNA polymerases move on copied DNA stochastically with forward/backward steps, proofreading, and context-dependent kinetics [8, 9, 10, 11, 12]. To capture non-ergodicity induced by compositional structure (e.g. AT- vs. GC-rich segments), we model replication as a block-diagonal Markov process on nucleotides.
Modeling stance.
[…]
@@
Let $\mathcal{X}=\{\mathrm{A},\mathrm{T},\mathrm{C},\mathrm{G}\}$ with two blocks $\mathcal{X}_{1}=\{\mathrm{A},\mathrm{T}\}$ and $\mathcal{X}_{2}=\{\mathrm{C},\mathrm{G}\}$. A one-step substitution kernel (copying) with block invariance (2) is
@@
Primitivity in this setting. For a two-state block [[1a,a],[b,1b]][[1-a,a],[b,1-b]], primitiveness is equivalent to $a>0$ and $b>0$ with $(a,b)\neq(1,1)$; in particular, $a,b\in(0,1)$ suffices. Under the mixture $\widetilde{T}=(1-\rho)T+\rho\mathcal{R}$, if $\mathcal{R}$ has strictly positive entries then any $\rho>0$ makes all effective entries positive, hence each block becomes primitive and $V(p_{n})\to 0$.
KLD potential with unique block invariants.
Assume that each block chain admits a unique invariant $\pi_{j}^{\ast}\in\mathcal{I}_{j}$. Then, combining (1) with (12) gives
$V(p_{n})\ =\ \sum_{j=1}^{2}w_{j}(p_{0})\,D_{\mathrm{KL}}\!\bigl(p_{n}^{(j)}\ \|\ \pi_{j}^{\ast}\bigr).$ (31)
By Theorem 1, $V$ is a Lyapunov function for the blockwise dynamics: $V(p_{n+1})\leq V(p_{n})$ for all $n$, and $V(p_{n})\to 0$ as $n\to\infty$.
Local detailed balance derivation.
Within each two-state block, write the undriven rates as
$A\xrightleftharpoons[\ \beta\ ]{\ \alpha\ }T\qquad\text{and}\qquad C\xrightleftharpoons[\ \delta\ ]{\ \gamma\ }G.$
At equilibrium (no chemical drive), detailed balance with the block’s stationary distribution $\pi^{\ast}$ gives
$\pi^{\ast}_{\!A}\,\alpha=\pi^{\ast}_{\!T}\,\beta\ \Rightarrow\ \frac{\alpha}{\beta}=\frac{\pi^{\ast}_{\!T}}{\pi^{\ast}_{\!A}},\qquad\frac{\gamma}{\delta}=\frac{\pi^{\ast}_{\!G}}{\pi^{\ast}_{\!C}}.$
When a proofreading cycle provides a chemical affinity $\Delta\mu>0$ per round-trip, the local detailed balance (cycle thermodynamics) shifts log rate odds by $\Delta\mu/k_{\mathrm{B}}T$:
$\ln\frac{\alpha^{\prime}}{\beta^{\prime}}=\ln\frac{\alpha}{\beta}+\frac{\Delta\mu}{k_{\mathrm{B}}T},\qquad\ln\frac{\gamma^{\prime}}{\delta^{\prime}}=\ln\frac{\gamma}{\delta}+\frac{\Delta\mu}{k_{\mathrm{B}}T}.$ (32)
Equivalently, in odds form,
$\frac{\alpha^{\prime}}{\beta^{\prime}}\;=\frac{\alpha}{\beta}\,\exp\!\Big(\frac{\Delta\mu}{k_{\mathrm{B}}T}\Big),\qquad\frac{\gamma^{\prime}}{\delta^{\prime}}\;=\frac{\gamma}{\delta}\,\exp\!\Big(\frac{\Delta\mu}{k_{\mathrm{B}}T}\Big).$ (33)
Thus, the drive fixes odds (ratios) but not individual rates; pinning $\alpha^{\prime},\beta^{\prime}$ (or $\gamma^{\prime},\delta^{\prime}$) separately requires an additional timescale constraint (for example, keeping $\alpha^{\prime}+\beta^{\prime}$ fixed), which we do not assume here. As is standard in stochastic thermodynamics, the exponential tilt of rate odds by the chemical cycle affinity follows from network cycle thermodynamics. Equations (32)–(33) encode local detailed balance for rate odds; they shape steady odds and relaxation speed but do not by themselves imply primitivity, which is a structural (irreducible+aperiodic) property; cf. Theorem 2.
Note. The exponential tilt in Eq. (33) biases odds but does not by itself specify the steady entropy production rate; the latter can remain positive even when $V\to 0$ under primitive blocks.
@@
In kinetic proofreading a chemical drive provides an affinity $\Delta\mu$ per cycle (e.g., via nucleotide hydrolysis). Within each invariant block, this drive tilts the two-state odds toward the proofreading-favored direction according to Eq. (33), sharpening the steady composition. Because $V$ decomposes over blocks, a stronger drive typically steepens the within-block basins (reducing the Bernoulli variance around $\pi_{j}^{\ast}$) and thus accelerates the monotone decay of $V$ guaranteed by Theorem 1. Crucially, the existence and monotonicity of $V$ do not require a drive; the drive only modifies the speed and the steady odds.
Primitivity of $2\times 2$ blocks and relation to Eq. (35).
For the two-state blocks in Example 2, a practical sufficient condition for block primitivity is strict positivity of every entry of the effective $2\times 2$ transition kernel: $\alpha_{\rm eff},\beta_{\rm eff},\gamma_{\rm eff},\delta_{\rm eff}\in(0,1)$. Then each block is irreducible and aperiodic (hence primitive), and (31) implies $V(p_{n})\to 0$. If $\alpha_{\rm eff}=0$ or $\beta_{\rm eff}=0$, irreducibility fails; if $\alpha_{\rm eff}=\beta_{\rm eff}=1$ (and analogously for CG), period-2 oscillations arise unless additional repair/proofreading steps restore positive diagonal entries. Equation (33) (Eq. 35) expresses how $\Delta\mu$ biases rate ratios; it determines the steady odds and relaxation speed but is not, by itself, a primitivity condition, which depends on positivity and aperiodicity of the effective kernels.
Notes. The sign of $\Delta\mu$ sets the favored direction; reversing the operating bias corresponds to $\Delta\mu\!\to\!-\Delta\mu$. At stationarity the block-wise odds inherit the exponential tilt: $\pi_{A}/\pi_{T}=\beta^{\prime}/\alpha^{\prime}$ and $\pi_{C}/\pi_{G}=\delta^{\prime}/\gamma^{\prime}$, consistent with the proofreading picture of Hopfield [13]. This connects proofreading accuracy to energetic cost in line with kinetic proofreading thermodynamics [13, 14].
@@
At each step, we record: $H(q_{n})$, $H_{\times}(n)=H(p_{n},q_{n})$, $D_{\mathrm{KL}}(p_{n}\|q_{n})$, and $V(p_{n})$, with monotone $V$ ensured by Theorem 1. Because $\widetilde{T}$ is block diagonal by construction, coarse transitions between $\{\mathrm{AT},\mathrm{GC}\}$ vanish at one step. Block-level confusion rates under small inter-block leakage are analyzed in the leakage section.
Readouts and figures.
The trajectories in Fig.2 show an increase in $H$ and $H_{\times}$, a decrease in $D_{\mathrm{KL}}(p_{n}\|q_{n})$, and a monotone decay of $V(p_{n})$, in agreement with Theorem 1. Figure 3 shows the potential landscape $V(p)=w_{1}V_{1}(x)+w_{2}V_{2}(y)$ on $(x,y)$, where $x$ is the A fraction in $\{\mathrm{A},\mathrm{T}\}$ and $y$ the C fraction in $\{\mathrm{C},\mathrm{G}\}$(Appendix D).
Figure 3: DNA block-diagonal replication: time series of informational metrics (nats).Shannon entropy, Cross entropy, KLD, and KLD potential. We iterate a four-state Markov chain on $\{\mathrm{A},\mathrm{T},\mathrm{C},\mathrm{G}\}$ with a block-diagonal kernel $\widetilde{T}=\mathrm{diag}(T_{1},T_{2})$ for $N=50$ steps. Base rates: $\alpha=0.020$, $\beta=0.010$, $\gamma=0.015$, $\delta=0.015$. Proofreading channel $\mathcal{R}$ with $\alpha^{\prime}=0.005$, $\beta^{\prime}=0.003$, $\gamma^{\prime}=\delta^{\prime}=0.004$ is mixed at $\rho=0.30$, giving effective rates $\alpha_{\!e}=0.0155$, $\beta_{\!e}=0.0079$, $\gamma_{\!e}=\delta_{\!e}=0.0117$ and block invariants $\pi_{1}^{\ast}=(0.3376,\,0.6624)$, $\pi_{2}^{\ast}=(0.5,\,0.5)$. Initial distribution $p_{0}=(0.6,0.1,0.2,0.1)$ so that $w_{1}(p_{0})=0.7$, $w_{2}(p_{0})=0.3$. At each step $q_{n}=p_{n}\widetilde{T}$ and we record: top-left, Shannon entropy $H(q_{n})$; top-right, cross-entropy $H(p_{n},q_{n})$; bottom-left, step divergence $D_{\mathrm{KL}}(p_{n}\|q_{n})$; bottom-right, KLD potential $V(p_{n})=w_{1}(p_{0})\,D_{\mathrm{KL}}(p_{n}^{(1)}\|\pi_{1}^{\ast})+w_{2}(p_{0})\,D_{\mathrm{KL}}(p_{n}^{(2)}\|\pi_{2}^{\ast})$. As replication proceeds, $H$ and $H_{\times}$ increase, while both $D_{\mathrm{KL}}(p_{n}\|q_{n})$ and $V(p_{n})$ decrease monotonically; the monotonicity of $V$ follows from Theorem 1.All metrics are reported in nats
The contour plot shows the KLD potential $V(p)=w_{1}D_{\mathrm{KL}}\!\bigl(p^{(1)}\!\parallel\!\pi_{1}^{\ast}\bigr)+w_{2}D_{\mathrm{KL}}\!\bigl(p^{(2)}\!\parallel\!\pi_{2}^{\ast}\bigr)$ on $(x,y)\in[0,1]^{2}$, where $x$ is the fraction A in the AT block and $y$ is the fraction C in the GC block. The unique minimum (blue marker) occurs at $(x^{\ast},y^{\ast})=(\pi_{1}^{\ast}(\mathrm{A}),\pi_{2}^{\ast}(\mathrm{C}))\approx(0.338,0.600)$, i.e., at the block-wise invariants determined by the effective rates. Because $V$ is a sum of blockwise KL terms, the level sets are nearly axis-aligned ellipses (here close to circular since $w_{1}=w_{2}=0.5$ and the curvatures are similar). A second-order expansion around $(x^{\ast},y^{\ast})$ gives $V(x,y)\approx\tfrac{w_{1}}{2\,\mu(1-\mu)}(x-\mu)^{2}+\tfrac{w_{2}}{2\,\nu(1-\nu)}(y-\nu)^{2}$ with $\mu=\pi_{1}^{\ast}(\mathrm{A})\approx 0.338$ and $\nu=\pi_{2}^{\ast}(\mathrm{C})=0.600$, yielding coefficients $\simeq 1.12$ and $\simeq 1.04$ (nats), respectively—consistent with the nearly isotropic contours. Increasing proofreading bias decreases $\mu(1-\mu)$ or $\nu(1-\nu)$, steepening the basin and accelerating the monotonic decay of $V(p_{n})$ (Theorem 1). The separable, convex landscape highlights that, under block invariance, the non-ergodic relaxation proceeds independently within blocks and converges to the reachable steady point $(x^{\ast},y^{\ast})$(Fig. 3).
Figure 4: KLD potential landscape $V(p)=w_{1}D_{\mathrm{KL}}(p^{(1)}\parallel\pi_{1}^{\ast})+w_{2}D_{\mathrm{KL}}(p^{(2)}\parallel\pi_{2}^{\ast})$ (nats) over $(x,y)\in[0,1]^{2}$, where $x$ is the A fraction in the AT block and $y$ is the C fraction in the GC block. Proofreading mixture $\rho=0.30$; AT rates $\alpha=0.020$, $\beta=0.010$, $\alpha^{\prime}=0.005$, $\beta^{\prime}=0.003$ give $\alpha_{\mathrm{eff}}=0.0155$, $\beta_{\mathrm{eff}}=0.0079$ and $\pi_{1}^{\ast}=(\beta_{\mathrm{eff}}/(\alpha_{\mathrm{eff}}+\beta_{\mathrm{eff}}),,\alpha_{\mathrm{eff}}/(\alpha_{\mathrm{eff}}+\beta_{\mathrm{eff}}))\approx(0.338,,0.662)$. GC rates (asymmetric) $\gamma=0.014$, $\delta=0.021$, $\gamma^{\prime}=0.004$, $\delta^{\prime}=0.006$ give $\gamma_{\mathrm{eff}}=0.0110$, $\delta_{\mathrm{eff}}=0.0165$ and $\pi_{2}^{\ast}=(\delta_{\mathrm{eff}}/(\gamma_{\mathrm{eff}}+\delta_{\mathrm{eff}}),,\gamma_{\mathrm{eff}}/(\gamma_{\mathrm{eff}}+\delta_{\mathrm{eff}}))\approx(0.60,,0.40)$. Block weights $w_{1}=w_{2}=0.5$. The blue marker indicates $(x^{\ast},y^{\ast})=(\pi_{1}^{\ast}(\mathrm{A}),,\pi_{2}^{\ast}(\mathrm{C}))\approx(0.34,,0.60)$. Grid $101\times 101$, colormap All values in nats.
Note on parameter sets. Figure 2 uses a symmetric GC block ($\gamma_{\!e}=\delta_{\!e}$; $\pi_{2}^{\ast}=(0.5,0.5)$), whereas Figure 3 uses an asymmetric GC block leading to $\pi_{2}^{\ast}=(0.60,0.40)$. The two figures illustrate, respectively, symmetric and biased within-block odds.
An information-theoretic second law
Standing assumptions.
We consider a discrete-time Markov evolution $p_{n+1}=Tp_{n}$ on a finite state space, with a block-invariant kernel $T=\mathrm{diag}(T_{1},\dots,T_{m})$ as in Sect. 1. Let $q_{n}:=p_{n}T$ denote the one-step image of $p_{n}$. The KLD potential is
$V(p)\;=\;\inf_{\pi\in\Pi(p_{0})}D_{\mathrm{KL}}(p\|\pi),$
where $\Pi(p_{0})$ is the reachable steady set (Def. (1)). By Theorem 1, $V(p_{n+1})\leq V(p_{n})$ for all $n$.
Step variables and their roles.
Define the potential drop and the one-step mismatch by
$\Delta V_{n}\;:=\;V(p_{n})-V(p_{n+1}),\qquad D_{n}\;:=\;D_{\mathrm{KL}}(p_{n}\|q_{n}),\quad q_{n}=p_{n}T.$ (34)
Nonnegativity holds as follows: (i) $\Delta V_{n}\geq 0$ by the Lyapunov monotonicity of $V$; (ii) $D_{n}\geq 0$ because KLD is nonnegative (and $D_{n}=0$ iff $p_{n}=q_{n}$).
Decomposition of the step production.
Define the dimensionless step production
$\mathcal{S}_{n}\;:=\;D_{n}+\Delta V_{n}.$ (35)
It gathers the irreversible loss within one update ($D_{n}$) and the progress toward the reachable steady set ($\Delta V_{n}$).
Proposition 1 (Nonnegativity and telescoping sum).
For any block-invariant trajectory $\{p_{n}\}_{n=0}^{N}$,
$\mathcal{S}_{n}\;\geq\;0\quad(n=0,1,\dots,N-1),\qquad\mathcal{S}^{[0:N)}\;:=\;\sum_{n=0}^{N-1}\mathcal{S}_{n}\;=\;\sum_{n=0}^{N-1}D_{n}\;+\;V(p_{0})-V(p_{N})\;\geq\;0.$ (36)
Proof.
By (34) and Theorem 1, $\Delta V_{n}\geq 0$; KLD nonnegativity gives $D_{n}\geq 0$, hence $\mathcal{S}_{n}\geq 0$. Summing yields $\sum_{n}\Delta V_{n}=V(p_{0})-V(p_{N})$. ∎
Equation (36) states that for any horizon $N$,
$\sum_{n=0}^{N-1}D_{\mathrm{KL}}(p_{n}\|p_{n}T)\;+\;V(p_{0})-V(p_{N})\;\geq\;0.$
Thus the accumulated one-step mismatch and the state-function decrease are jointly nonnegative—an information-theoretic second-law inequality under non-ergodic (block-invariant) dynamics. The key departure from the classical ergodic setting is that $V$ references the reachable steady set rather than a single invariant distribution.
Theorem 2 (Primitivity of block kernels).
Let $T$ be a block transition kernel. If each block is primitive (irreducible and aperiodic), then the system admits a unique invariant distribution within each block, and the dynamics converge to this invariant.
Theorem 3 (Second law for non-ergodic replication: minimal form).
Equivalently to Proposition 1, for all $n$ and $N\geq 1$,
$\mathcal{S}_{n}\;\geq\;0,\qquad\mathcal{S}^{[0:N)}\;\geq\;0.$ (37)
In physical units, set $\sigma_{\rm rep}^{(n)}:=k_{\mathrm{B}}T\,\mathcal{S}_{n}$ and $\Sigma_{\rm rep}^{(N)}:=k_{\mathrm{B}}T\,\mathcal{S}^{[0:N)}$.
Proof.
Immediate from Proposition 1. ∎
Remark 1 (Tightness conditions).
$\mathcal{S}_{n}=0$ iff $D_{n}=\Delta V_{n}=0$, i.e., (i) $p_{n}=q_{n}$ (a fixed point of $T$), and (ii) $V(p_{n+1})=V(p_{n})$, which requires blockwise tightness of DPI as in Remark 1. Otherwise, $\mathcal{S}_{n}>0$.
Thermodynamic reading.
$V$ measures the distance to the reachable steady set; its drop $k_{\mathrm{B}}T\,\Delta V_{n}$ plays the role of an informational free-energy decrease, while $k_{\mathrm{B}}T\,D_{n}$ quantifies stepwise dissipative loss of distinguishability. Hence $\mathcal{S}_{n}$ acts as a nonnegative production:
$\text{(instantaneous dissipation)}\ +\ \text{(state-function decrease)}\;\geq\;0.$
The monotonicity of $V$ uses Lemma 2 and DPI applied blockwise, $D_{\mathrm{KL}}(T_{j}r\|\pi_{j})\leq D_{\mathrm{KL}}(r\|\pi_{j})$ with $\pi_{j}T_{j}=\pi_{j}$. If blocks leak, $\Pi(p_{0})$ ceases to align with the one-step map and $V$ may increase in one step We describe the blocks leak in Appendix A.
Extensions (time-dependent/continuous-time).
(1) Time-dependent kernels. If each $T_{n}$ shares the same block partition holds blockwise, then $V(p_{n+1})\leq V(p_{n})$ still holds and the theorem remains valid with $q_{n}=p_{n}T_{n}$. (2) Continuous-time CTMC. For generator $Q$ and propagator $K(h)=e^{hQ}$, small-$h$ expansions give $D_{\mathrm{KL}}(p\|pK(h))=\tfrac{h^{2}}{2}\,\dot{\mathcal{I}}(p;Q)+O(h^{3})\geq 0$, while $V(p(t+h))\leq V(p(t))$ for all $h>0$ (Theorem 2 ensures positivity/primitivity). Integrating yields the continuous-time analogue: